what is the sum of the measurements of the interior angles of a 28-gon?

I'm trying to help my son with his homework and help would be great!

Answer by checkley71(8403)

A 28-GON HAS 28 INTERIOR ANGLES.
THE SUM OF INTERIOR ANGLES OF AN X-GON=180(X-2)
180(28-2)=180*26=4,680 DEGREES IS THE ANSWER.